OpenStudy (anonymous):

Why can a subtraction problem be rewritten as an addition problem without changing the outcome? Give a real-world scenario where this might be modeled.

OpenStudy (anonymous):

Because the definition of subtraction in the integers is the addition of the inverse. A.k.a:\[a-b=a+-b\]Where\[a, b \in \mathbb{Z}\]And, where\[-b\]is defined as the solution k to:\[k+b=0\] A real world problem? I walked 5 steps backwards and took three steps forwards, what is my outcome?